OpenStudy (anonymous):

what is the derivative of cosxsin3x?

OpenStudy (anonymous):

\[(-sinx)(\sin3x)+cosx(\cos3x)(3)\] \[-\sin^23x+3\cos^23x\]

OpenStudy (zarkon):

@winterfez you can't combing the sines and cosines like that

OpenStudy (zarkon):

*combine

OpenStudy (anonymous):

is cosx time sin3x, thus I can use addition for derivative

OpenStudy (zarkon):

\[(-\sin(x))\sin(3x)\ne -\sin^2(3x)\]

OpenStudy (anonymous):

oh that was what you were pointing out..THANKS..didnt see that mistake